Mellie is conducting a test on mold spores on bread. She uses 6 slices of bread to compare two strains of mold. She applies one strain to the left side of the bread and one strain to the right side. She flips a coin to decide which strain goes on the right side of the bread. The mold spores that appear on each side are counted and she records them in a table.

Bread Number of Spots for Strain 1 Number of Spots for Strain 2

1

30

19

2

21

17

3

15

15

4

13

12

5

6

9

6

8

4



If Mellie is to perform an appropriate t-test to determine if there is a difference in the mean number of spots per piece of bread produced by the two strains, how many degrees of freedom should she use?

4

5

6

10

12
Answer #1

If Mellie is to perform an appropriate t-test to determine if there is a difference in the mean number of spots per piece of bread produced by the two strains.

sample size 1 ( n1) = 6

sample size 2 ( n2 ) = 6

If we assume the population variance are equal then degrees of freedom are,

n1 + n2 - 2 = 6 + 6 - 2 = 10

Degrees of freedom = 10

Note: if we do not assume population standard deviations are equal then

degrees of freedom = min ( 6 - 1, 6 - 1) = 5